# Hire Your Anthropologist! Rethinking Culture Benchmarks Through an Anthropological Lens

### v1

@@ 3.1 Mapping Benchmarks Onto the Taxonomy
We review 20 prominent cultural benchmarks in NLP and map them to the four cultural dimensions. Figure 2 presents the category breakdown, with full descriptions in Table 1. Each of the four parts of our framework highlights a distinct facet of cultural life and can fit different evaluation goals. However, most of the prominent NLP benchmarks to date lean on two narrow frameworks: culture as knowledge and culture as preference, while treating culture as a set of static facts or simple surveyable attitudes. In contrast, anthropological work views culture as dynamic, situated, and relational. In our recommendations (§5), we will expand on how to use this taxonomy in future work.
@@ 4 Methodological Limitations and Ways Forward
This section identifies six methodological limitations that flatten cultural complexity in current NLP benchmarks. We argue that because of these methodological choices, the concept of culture is often unidimensional and flattened they are: (a) platform bias favoring Western, urban demographics; (b) conflating nation-states with cultural boundaries; (c) treating individual annotators as cultural representatives; (d) reducing moral reasoning to survey scales; (e) assuming cultural consensus where disagreement exists; and (f) stripping context from cultural scenarios. These practices systematically undermine all four dimensions of our taxonomy, knowledge, preference, dynamics, and bias, by imposing static, essentialist frameworks that contradict anthropological understandings of culture as contested, situated, and fluid. Table 1 maps these limitations across benchmarks, revealing how methodological choices constrain conceptual possibilities and pointing to the need for more reflexive, context-sensitive, and theoretically grounded evaluation methods.
@@
Table 1: Presence of Methodological Limitations Across 20 Cultural Benchmarks. Each column indicates the presence (✓) or absence (–) of a limitation defined in Section 4: Platform Bias4.1), Nation-State4.2), Annotation4.3), Moral Simplification4.4), Consensus4.5), Lack of Context4.6).
@@ 4.1 Platform Bias and Demographic Skew
By sampling from narrow, digitally privileged demographics, platform bias undermines all four dimensions of our taxonomy (knowledge, preference, dynamics, bias). Many cultural and bias benchmarks draw from Reddit, Twitter, or Wikipedia, which mirror Western, male, urban, and digitally literate users. Social Chemistry 101 inherits Reddit’s community biases, and datasets like BBQ, BOLD, CulturalAtlas, and others import similar skews; the effect extends to BLEND, CASA, CDEval, CulturalBench, LLM-GLOBE, SeaExam, and X-FACTR. Treating such homogeneous pools as globally representative risks erasing offline, indigenous, rural, or otherwise marginalized worldviews.
Partial remedies exist. FLEAD (Federated Learning to Exploit Annotator Disagreements) Rodríguez-Barroso et al. (2024) models each annotator’s perspective to surface minority viewpoints that majority aggregation can suppress, yet its reach is limited by sparse annotator metadata and the exclusion of communities without digital access. More promising directions treat culture as situated and temporal, emphasizing multiplicity over universality. BLEND Myung et al. (2025) advances this by evaluating culturally situated knowledge across 16 regions and 13 languages using contemporary community-informed data, and SEACrowd Lovenia et al. (2024) shows what large-scale community collaboration can achieve with over 400 Southeast Asian contributors.
@@ 4.2 Nation-State as a Proxy for Culture
Several NLP benchmarks adopt nationality as the primary cultural frame, assuming country boundaries map to homogeneous cultures. SeaExam and CulturalAtlas assign norms by country with statements like “In Chinese culture…,” while BLEND, CDEval, LLM-GLOBE, and X-FACTR rely on national groupings that obscure internal ethnic, linguistic, and religious differences. SeaExam aggregates Singapore, Malaysia, and Thailand under one label, and LLM-GLOBE conflates language with culture (for example, Arabic as “Middle Eastern culture”), overlooking diasporic contexts.
Social science shows why this is problematic. Gupta and Ferguson (1992) argues that nation states are political constructs shaped by transnational flows and historical contingencies, not reliable proxies for cultural coherence. BLEND demonstrates progress by combining intra- and cross-national perspectives, and Jiraibench Xiao et al. (2025) probes how meanings shift across discourse types, platforms, and diasporic communities. Jiraibench is an exploratory first step; sustained research is needed to develop richer, more situated evaluations that reflect contested, plural, and evolving cultural life.
@@ 4.3 Representation Fallacy Through Annotation
Many benchmarks assume that a few annotators can represent the cultural norms of a group, affecting BBQ, BLEND, CASA, CDEval, SEAExam, LLM-GLOBE and others. SEAExam collapses Singapore, Malaysia, and Thailand into one label, and LLM-GLOBE rarely reports annotator demographics. Treating such labels as comprehensive removes the variation by class, gender, generation, and religion. Using LLMs as “cultural judges” without external validation amplifies the problem. As Abu-Lughod (1991) and Mahmood (2005) note, representation is not only who speaks, but also how voices are situated within power relations. The result is a flattening of Culture-as-Knowledge into generic summaries and Culture-as-Preference into unexamined averages.
@@ 4.4 Moral Simplification through Survey Format Overload
Many “cultural reasoning” benchmarks reduce Culture-as-Preference and Culture-as-Dynamics to static survey responses, collapsing nuanced judgments into Likert scales or binaries. CASA, CDEval, Civics, Moral Bench, Moral Stories, NormAd, and Social Chemistry 101 rely on simplified prompts that flatten context. Moral Bench, grounded in Moral Foundations Theory built from WEIRD samples, treats ethics as consensus scores and projects a narrow frame as universal. Anthropological work shows why this fails: moral life is context-dependent, narratively situated, and historically contingent, not a set of fixed rules Zigon (2007); Laidlaw (2013).
Hybrid approaches such as Social Bias Frames Sap et al. (2020) replace surveys with scenario-based tasks that present culturally contingent dilemmas. These reveal value conflicts flattened in traditional formats, though scenarios often reflect Western paradigms of autonomy rather than collectivist frameworks Hofstede (1980). Progress requires preferring open-ended tasks that solicit narrative explanations alongside ratings.
@@ 4.5 Assumption of Cultural Consensus
Many benchmarks treat aggregated judgments as correct answers, assuming cultural agreement on norms,which eliminates the contested nature essential to Culture-as-Dynamics. CASA, CDEval, Moral Bench, Moral Stories, NormAd, and Social Chemistry 101 use majority votes as ground truth. Social Chemistry 101 collects explanations showing diverse reasoning, but still trains models on most common answers. This treats disagreement as noise, rather than recognizing culture as a site of ongoing negotiation. As Gal (2015) and Tsing (2005) emphasize, disagreement is a fundamental feature of cultural life, not a flaw.
@@ 5 Recommendations
A more multidimensional understanding of culture can emerge by addressing these methodological shortcomings, such as supplementing national and demographic proxies with data from local communities, capturing dynamic practices through real-life scenarios and narratives, and involving annotators from diverse backgrounds to reflect internal variation.
@@ 5.1 Social Science Collaboration
An example of this approach is AlKhamissi et al., who developed interdisciplinary annotation protocols and evaluation frameworks by establishing sustained collaboration between computer scientists and social scientists from the inception of the project through analysis, illustrating both challenges and best practices to integrate anthropological insights into the design of the NLP benchmark. This work, along with surveys such as NLP for Social Good Karamolegkou et al. (2025), Artz’s “Anthropology and AI: A Framework for Mutually Beneficial Collaboration”Artz (2025), and initiatives by Azimuth Labs, underscore that a deeply embedded social science partnership leads to richer, more robust and culturally attuned AI systems.
@@ 5.4 Critical Annotation
A critical annotation approach must move beyond simply measuring annotator disagreement, recognizing that annotation itself is a social and interpretive act deeply shaped by the power relations and positionality within annotation work. While existing tools for calculating inter-annotator agreement help quantify variation, this is only a partial solution: a single annotator’s choices reflect not the entirety of a culture or religion, but their own relationship with and perspective on those traditions—a performance influenced by social, economic, and institutional contexts. Drawing from anthropology’s tradition of ’writing culture’ Clifford and Marcus (1986), critical annotation recommends protocols that allow ambiguity and narrative, incorporating continuous scales, narrative storytelling, and space for dissent and self-reflection rather than forcing binary consensus. Methods should encourage iterative, community-driven processes where annotation guidelines and ethical dilemmas are refined collaboratively, capturing multiple, contested meanings and the complexity of lived experience.
@@ 5.5 Mapping Controversy
Tools should be developed to systematically map zones of controversy in datasets, institutionalizing plural ’ground truths’ where contestation and disagreement are recognized as valuable data signals. Rather than erasing dissent, benchmarks should record, analyze and represent sites of active dispute as key sources of interpretive richness. Empowering communities to identify which disagreements matter and leveraging critical annotation practices shifts evaluation away from artificial consensus and allows AI systems to better reflect the multiplicity and contested nature of culture.
Examples of controversy mapping and plural ground truth approaches can be found in interdisciplinary algorithm studies and recent NLP benchmarks that explicitly document disagreement. Munk et al. have mapped algorithmic controversies in the scientific literature by tracking how different communities contest and reinterpret algorithmic outputs, making dissent part of the data landscape. In NLP, FLEAD’s consensus building protocol preserves annotator differences, and projects like Social Chemistry 101 retain gradient judgments to reflect multiple interpretations of social norms rather than flattening to a single answer.
@@ 5.6 Using the Full Taxonomy
Our taxonomy serves as an integrative framework that bridges persistent disconnects between different conceptualizations of culture used in NLP, evaluation, and anthropological theory. By systematically distinguishing and relating culture as knowledge, culture as preference, culture as dynamics, and culture as bias, it illuminates both the methodological fragmentation and the scope for meaningful synthesis in future benchmarks. One of the main limitations of existing resources is their tendency to focus narrowly on a single dimension, which risks essentializing culture (by treating it as static knowledge), universalizing group values, or reducing sociocultural dynamics to binary bias detection. As a critique, this compartmentalization overlooks the interplay between what is known, what is valued, how meaning is enacted, and how stereotypes or harms emerge in context. Moving forward, benchmark design should more deliberately combine elements from across this taxonomy, rather than isolating them, in order to create richer, more contextually aware, and theoretically grounded resources. Such an approach would not only support interdisciplinary research, but would also encourage a more dynamic, intersectional, and nuanced evaluation of the relationship of NLP models with culture.
@@ 6 Conclusion
Culture is not a static set of facts to memorize; it is a living and contested practice negotiated daily. Current NLP benchmarks often reduce culture to trivia, so fixes like diverse annotators, localized fine-tuning, and disagreement-aware metrics capture only fragments. Drawing on anthropology, we propose a flexible taxonomy and an interdisciplinary pipeline where social scientists collaborate with NLP researchers across design, data, modeling, and evaluation. Social theory can guide scenario-based, context-rich prompts and treat annotator disagreement as a signal, while NLP methods scale these insights. Sustained collaboration replaces ad hoc work with iterative co-design and joint interpretation, yielding benchmarks that reflect fluid identities and models that engage culture as lived practice, improving both bias mitigation and interpretability.
@@
Our study focuses primarily on benchmarking methodologies rather than assessing the real-world impact of cultural benchmarks on LLM applications. Although improving benchmarks is a crucial step toward cultural sensitivity in AI, future research should investigate how these advancements translate to tangible improvements in AI-human interactions across various domains, including education, governance, healthcare, and cross-cultural communication. Without real-world deployment analysis, the long-term effectiveness of improved cultural benchmarks remains an open question.
In addition, our position paper focuses on text-based evaluations, omitting vision-oriented culture benchmarks that interrogate the understanding of models of imagery, symbolism, and geospecific visual practices. Multimodal datasets such as CulturalVQA (visual question answering on food, clothing and rituals drawn from 11 countries) Nayak et al. (2024), CVQA Romero et al. (2024) (a culturally diverse and multilingual VQA set spanning 28 nations and 26 languages), and WorldCuisine Winata et al. (2025) (a dataset of world cuisines) demand analytical framework (e.g., visual-semantics grounding, region-specific iconography, and image–text alignment biases) that differ fundamentally from the linguistic constructs underlying our framework. Incorporating them would therefore require distinct taxonomies and fairness diagnostics, which we flag as an important avenue for future multimodal culture work.
We also recognize that we are not the first to introduce a taxonomy of cultural understanding in NLP Liu et al. (2025b). However, in contrast, our four-part taxonomy distinguishes between culture-as-knowledge (factual/symbolic information), culture-as-preference (shared values and moral attitudes), culture-as-dynamics (performative, enacted, context-dependent practices), and culture-as-bias (patterns of stereotyping and harm reduction). Our framework is not only descriptive, mapping how culture is conceptualized, but also prescriptive, challenging single-axis approaches, and recommending that future benchmarks intentionally combine these dimensions to better capture the complexity and contested nature of cultural life.
@@
Appendix A History of Culture
From traits to contexts.
Early anthropology, most clearly in Boas’ cultural relativism, rejected universal rankings of peoples and argued that cultural phenomena must be read in their historical and local contexts (Boas, 1911). This moved the field away from trait lists to contextual explanations.
Interpretive turn.
Geertz recast culture as ’webs of significance’, urging a thick description of symbols, rituals, and everyday acts to understand meaningmaking (Geertz, 1973). This positioned interpretation, not only measurement, at the center of cultural analysis.
Writing culture.
The “writing culture” turn highlighted that accounts of culture are constructed through the researcher’s situated narrative, not simply discovered “out there” (Clifford and Marcus, 1986). Reflexivity and positionality became methodological commitments.
Performance and interaction.
Goffman’s lens treated culture as enacted through roles, audience design, and situational constraints (Goffman, 1959). Cultural competence is thus interactional and contingent, a view that motivates our attention to context sensitive evaluation in the main text.
Appendix B Taxonomy Details and Derivation
This appendix explains how we derived, applied, and validated the four-part taxonomy for cultural evaluation in NLP. The goal is to make our coding process transparent while keeping the main text focused. We describe the corpus and versioning we analyzed, the coder workflow and decision rules we used, and the operational cues that anchored each lens.
B.1 Derivation Procedure
We compiled a corpus of 20 cultural benchmarks and closely related resources that are widely cited in recent NLP work. For each resource, we recorded version identifiers, access dates, and any published annotation or scoring guidelines. Two coders independently reviewed task definitions, item formats, instructions to annotators, and evaluation metrics. Each coder proposed lens assignments with short rationales that referenced concrete aspects of item design or scoring. We then held an adjudication pass to resolve disagreements, documented the rationale for the final labels, and logged any unresolved ambiguities for follow-up sensitivity checks. This procedure emphasized stable signals in the written protocols rather than impressions from a small number of example items.
B.2 Coding Scheme and Decision Rules
Each benchmark received a primary lens that reflected the central evaluation target and, when warranted, a secondary lens to acknowledge additional design dependencies. We allowed multiple lenses only when both the item format and the scoring rubric would be materially altered by removing that dimension. When knowledge and preference co-occurred, the use of acceptability ratings, Likert scales, or aggregation to a group label was taken as evidence that preference should be primary. When a benchmark scored role-sensitive behavior, register selection, mitigation, or adaptation to context, we labeled dynamics as primary. Bias was labeled primary when the stated goal was to measure harm, disparity, or stereotype reproduction across groups, regardless of whether cultural facts or moral ratings appeared as supporting context. All decisions were based on observable artifacts, such as prompt templates, instructions, and metric definitions, rather than inferences about intended use.
B.3 Operational Cues for Each Lens
Culture-as-knowledge is indicated when items require factual recall or recognition tied to cultural artifacts, calendars, practices, or named entities that can be verified against grounded sources. Typical scoring relies on exact match, span match, or objective keying, and correctness does not depend on the social situation in which the answer is given.
Culture-as-preference is indicated when items elicit attitudes, norms, or moral judgments and when evaluation aggregates individual responses into a shared profile. Common signals include Likert style ratings, pairwise moral or social acceptability judgments, and scoring keyed to alignment with an aggregate label rather than to a single verifiable fact. Explanations may be collected, but the metric treats the aggregate judgment as the target.
Culture-as-dynamics is indicated when items are situated scenes or dialogues where roles, relationships, stakes, and settings shape what counts as appropriate language use. Here, success depends on interactional choices such as register selection, mitigation strategies, face management, and norm negotiation. Scoring rubrics explicitly reference contextual appropriateness or turn design, rather than only the final outcome.
Culture-as-bias is indicated when the central objective of the benchmark is to measure harms or disparities related to identity. Typical signals include prompts designed to surface stereotype triggers, metrics that report group gaps or bias directionality, and protocols that treat harm detection and reduction as the primary objective. The presence of cultural facts or moral ratings in such datasets serves context but does not redefine the evaluation target.
Appendix C Description of Reviewed Benchmarks

### v2

@@ 4.2 Mapping Benchmarks Onto the Taxonomy
We review 20 prominent cultural benchmarks in NLP and map them to the four cultural dimensions. Datasets such as CultureAtlas Feng (2024), BLEnD Myung et al. (2025), and CulturalBench Chiu et al. (2024) exemplify Culture-as-Knowledge, focusing on factual recall tied to national or regional identity. Culture-as-Preference is illustrated by benchmarks such as Social Chemistry 101 Forbes et al. (2020), MoralBench Ji et al. (2024), and NormAd Rao et al. (2025), where models are evaluated for alignment with aggregated moral or social judgments. Culture-as-Dynamics appears in datasets such as CoCo-LM Meng et al. (2021), BLEnD, and MoralStories, foregrounding the context-sensitive nature of language and emotion. Finally, Culture-as-Bias is captured by benchmarks such as BBQ Parrish et al. (2022), Social Bias Frames Sap et al. (2020), and BOLD Dhamala et al. (2021), which examine how models reproduce identity-based harms.
Figure 2 presents the category breakdown. Each of the four parts of our framework highlights a distinct facet of cultural life and can fit different evaluation goals. However, most of the prominent NLP benchmarks to date lean on two narrow frameworks: culture-as-knowledge and culture-as-preference, while treating culture as a set of static facts or simple surveyable attitudes. In contrast, anthropological work views culture as dynamic, situated, and relational. In our recommendations (§6), we will expand on how to use this taxonomy in future work.
@@ 5 Methodological Limitations and Ways Forward
This section identifies six methodological limitations that flatten cultural complexity in current NLP benchmarks. We argue that because of these methodological choices, the concept of culture is often unidimensional and flattened.
The limitations are: (a) Platform bias favoring Western, urban demographics; (b) Conflating nation states with cultural boundaries; (c) Treating individual annotators as cultural representatives; (d) Reducing moral reasoning to survey scales; (e) Assuming cultural consensus where disagreement exists; and (f) Stripping context from cultural scenarios.
These practices systematically undermine all four dimensions of our taxonomy: knowledge, preference, dynamics, and bias, by imposing essentialist frameworks that contradict complex understandings of culture as contested, situated, and fluid. We also show where benchmarks do a good job in finding creative ways around these methodological limitations. Table 1 maps these limitations across benchmarks, revealing how methodological choices constrain conceptual possibilities and pointing to the need for more reflexive, context-sensitive, and theoretically grounded evaluation methods.
@@
Table 1: Presence of Methodological Limitations Across 20 Cultural Benchmarks. Each column indicates the presence (✓) or absence (–) of a limitation defined in Section 5: Platform Bias5.1), Nation-State5.2), Annotation5.3), Moral Simplification5.4), Consensus5.5), Lack of Context5.6).
@@ 5.1 Platform Bias and Demographic Skew
Many cultural benchmarks are drawn from Reddit, X, or Wikipedia, which mostly mirror Western, male, urban, and digitally literate users. This privileges certain demographics over others. Social Chemistry 101 inherits Reddit’s community biases, and datasets like BBQ, BOLD, CulturalAtlas, and others import similar skews; the effect extends to CASA, CDEval, CulturalBench, LLM-GLOBE, SeaExam, and X-FACTR. Treating such homogeneous pools as globally representative risks erasing offline, indigenous, rural, or otherwise marginalized worldviews.
Digital anthropology demonstrates that digitally mediated expressions, such as posts and opinions on social networks, do not automatically reflect the full complexity of ’real social facts’, as recognized in ethnographic research Miller (2018). Miller stresses that digital platforms produce distinctive forms of sociality and cultural interaction, deeply shaped by the technological frameworks and demographics of their users; Interpreting these online traces as direct representations of offline life risks missing the broader social, cultural, and material contexts in which meaning is formed Horst and Miller (2012); Miller (2018). Anthropological analysis insists on holistic ethnography, situating digital worlds within the complexities of larger social relations, and cautioning against reducing the mediation of online experience to simple, universal truths Horst and Miller (2012); Miller (2018). Digital is not a transparent lens but an active mediator: sometimes intensifying difference, sometimes shaping whose voices become visible or invisible, and always requiring careful contextual understanding to avoid misrepresenting the realities of diverse communities Horst and Miller (2012); Miller (2018).
Partial remedies exist. FLEAD (Federated Learning to Exploit Annotator Disagreements) (Rodríguez-Barroso et al., 2024) models each annotator’s perspective to surface minority viewpoints that majority aggregation can suppress, yet its reach is limited by the sparse annotator metadata and the exclusion of communities without digital access.
@@ 5.2 Nation-State as a Proxy for Culture
Several NLP benchmarks adopt nationality as the primary cultural framework, assuming that country boundaries are mapable to homogeneous cultures. SeaExam and CulturalAtlas assign norms by country with statements such as ’In Chinese culture…’, while BLEND, CDEval, LLM-GLOBE and X-FACTR rely on national groups that obscure internal ethnic, linguistic and religious differences. SeaExam aggregates Singapore, Malaysia, and Thailand under one label, and LLM-GLOBE conflates language with culture (for example, Arabic as “Middle Eastern culture”), overlooking diasporic contexts. In agreement with our work,Havaldar et al. also shows how, instead of defining cultures, LLMs instead use proxies using nation states.
Social science scholarship underscores why such assumptions are fundamentally problematic. Gupta and Ferguson argues that nation states are politically constructed entities, shaped by transnational flows, colonization histories, and shifting political economies, rather than natural or coherent cultural units. Treating national borders as cultural boundaries flattens complex identities and histories, ignoring the dynamic, overlapping, and contested nature of culture that ethnographic research reveals. This critique aligns with broader calls within anthropology and cultural studies to move beyond the nation state as the default unit of cultural analysis and instead to address the multiplicity and fluidity of cultural formations Gupta and Ferguson (1992); Appadurai (1996).
BLEND demonstrates progress by combining intra- and cross-national perspectives, and Jiraibench (Xiao et al., 2025a) investigates how meanings change between discourse types, platforms, and diasporic communities. Jiraibench is an exploratory first step; sustained research is needed to develop richer, more situated evaluations that reflect contested, plural, and evolving cultural life.
@@ 5.3 Representation Fallacy Through Annotation
Many benchmarks assume that a few annotators can represent the cultural norms of a group. This affects BBQ, BLEND, CASA, CDEval, SEAExam, LLM-GLOBE and others. SEAExam collapses Singapore, Malaysia, and Thailand into one label, and LLM-GLOBE rarely reports annotator demographics. Treating such labels as comprehensive removes the variation by class, gender, generation, and religion. Using LLMs as “cultural judges” without external validation amplifies the problem.
As Abu-Lughod (1991) and Mahmood (2005) note, representation is not only who speaks, but also how voices are situated within power relations. Little is understood of how annotation choices, sampling practices, or positionality of annotators shape the contours of what counts as ’culture’ in these data sets. This methodological opacity risks reinforcing normative assumptions, perpetuating a narrow view of cultural diversity that can marginalize less visible voices and mask tensions within communities. Without reflexive attention to which experiences are foregrounded and which are silenced, benchmarks may inadvertently produce a shallow or exclusionary portrait of cultural life.
@@ 5.4 Moral Simplification through Survey Format Overload
Many “cultural reasoning” benchmarks reduce Culture-as-Preference and Culture-as-Dynamics to survey responses, collapsing nuanced judgments into Likert scales or binaries. CASA, CDEval, Civics, Moral Bench, Moral Stories, NormAd, and Social Chemistry 101 rely on simplified prompts that flatten context. Moral Bench, grounded in Moral Foundations Theory built from WEIRD samples, treats ethics as consensus scores and projects a narrow frame as universal. Anthropological work shows why this fails: moral life is context-dependent, narratively situated, and historically contingent, not a set of fixed rules (Zigon, 2007; Laidlaw, 2013).
Hybrid approaches such as Social Bias Frames Sap et al. (2020) replace surveys with scenario-based tasks that present culturally contingent dilemmas. These reveal value conflicts flattened in traditional formats, though scenarios often reflect Western paradigms of autonomy rather than collectivist frameworks (Hofstede, 1980). Progress requires preferring open-ended tasks that solicit narrative explanations alongside ratings.
@@ 5.5 Assumption of Cultural Consensus
Many benchmarks treat aggregated judgments as correct answers, assuming cultural agreement on norms, which eliminates the contested nature essential to Culture-as-Dynamics. CASA, CDEval, Moral Bench, Moral Stories, NormAd, and Social Chemistry 101 use majority votes as ground truth. Social Chemistry 101 collects explanations showing diverse reasoning, but still trains models on the most common answers. This treats disagreement as noise, rather than recognizing culture as a site of ongoing negotiation. As Gal (2015) and Tsing (2005) emphasize, disagreement is a fundamental feature of cultural life, not a flaw.
@@
More promising directions treat culture as situated and temporal, emphasizing multiplicity over universality. BLEND Myung et al. (2025) advances this by evaluating culturally situated knowledge in 16 regions and 13 languages using contemporary community-informed data, and SEACrowd Lovenia et al. (2024) shows what large-scale community collaboration can achieve with more than 400 Southeast Asian contributors.
@@ 5.6 Decontextualized Prompts and Abstracted Norms
Blodgett et al. (2024) argue the abstraction of culture in benchmark design preserves the illusion of neutrality, but erases contestation and dissent, privileging dominant perspectives and rendering marginal ones invisible. The most promising anthropological approach moves evaluation toward observing model behaviors within context-rich simulations and attending to gaps between stated outputs and embedded decisions. Recent work demonstrates that models can exhibit significant disparities along sociodemographic lines in complex settings, even when explicit bias is minimal.
@@ 6 Recommendations
These six design choices arise from the drive for scalability and quantifiability of NLP, which conflicts with the anthropological notions of situated practice. Here is our list of recommendations for ways forward.
@@ 6.1 Social Science Collaboration
An example of this approach is AlKhamissi et al., who developed interdisciplinary annotation protocols and evaluation frameworks by establishing sustained collaboration between computer scientists and social scientists from the inception of the project through analysis, illustrating both challenges and best practices to integrate anthropological insights into the design of the NLP benchmark. This work, along with surveys such as NLP for Social Good (Karamolegkou et al., 2025), Anthropology and AI: A Framework for Mutually Beneficial Collaboration (Artz, 2025), and other initiatives, underscore that a deeply embedded social science partnership leads to richer, more robust and culturally attuned AI systems.
@@ 6.4 Critical Annotation
A critical annotation approach must move beyond simply measuring annotator disagreement, recognizing that annotation itself is a social and interpretive act deeply shaped by the power relations and positionality within annotation work. While existing tools for calculating inter-annotator agreement help quantify variation, this is only a partial solution: a single annotator’s choices reflect not the entirety of a culture or religion, but their own relationship with and perspective on those traditions—a performance influenced by social, economic, and institutional contexts. Drawing from anthropology’s tradition of ’writing culture’ (Clifford and Marcus, 1986), critical annotation recommends protocols that allow ambiguity and narrative, incorporating continuous scales, narrative storytelling, and space for dissent and self-reflection rather than forcing binary consensus. Methods should encourage iterative, community-driven processes where annotation guidelines and ethical dilemmas are refined collaboratively, capturing multiple, contested meanings and the complexity of lived experience.
@@ 6.5 Mapping Controversy
Tools should be developed to systematically map zones of controversy in datasets, institutionalizing plural ’ground truths’ where contestation and disagreement are recognized as valuable data signals. Rather than erasing dissent, benchmarks should record, analyze, and represent sites of active dispute as key sources of interpretive richness. Empowering communities to identify which disagreements matter and leveraging critical annotation practices shifts evaluation away from artificial consensus and allows AI systems to better reflect the multiplicity and contested nature of culture.
Examples of controversy mapping and plural ground truth approaches can be found in interdisciplinary algorithm studies and recent NLP benchmarks that explicitly document disagreement. Munk et al. (2024) have mapped algorithmic controversies in the scientific literature by tracking how different communities contest and reinterpret algorithmic outputs, making dissent part of the data landscape. In NLP, FLEAD’s consensus-building protocol preserves annotator differences, and projects like Social Chemistry 101 retain gradient judgments to reflect multiple interpretations of social norms rather than flattening to a single answer.
@@ 6.6 Using the Full Taxonomy
Our taxonomy serves as an integrative framework that bridges persistent disconnects between the diverse conceptualizations of culture found in NLP research, evaluation design, and anthropological theory. By systematically distinguishing and relating culture-as-knowledge, culture-as-preference, culture-as-dynamics, and culture-as-bias, it reveals both methodological fragmentation and the potential for meaningful synthesis in future benchmarks. We recommend using the full taxonomy of culture when thinking about how to approach the question of culture within NLP. We also encourage researchers not only to use all four dimensions of this taxonomy when conceptualizing culture in NLP, but also to place it in conversation with other major frameworks: Adilazuarda et al. on demographic and semantic proxies; Liu et al. on ideational, linguistic and social dimensions; and Pawar et al. on multimodal and interactional approaches to cultural awareness. When used together, these frameworks can advance both theoretical and evaluative sophistication, linking conceptual precision with methodological design. Moving forward, benchmark development should integrate all aspects of our taxonomy while engaging these complementary models to produce more dynamic, intersectional, and contextually grounded evaluations of how NLP systems represent and negotiate culture.
@@ 7 Conclusion
Current NLP benchmarks often reduce culture to trivia, to fixes such as diverse annotators, localized fine-tuning, and disagreement-aware metrics capture only fragments. Drawing on anthropology, we propose a flexible taxonomy and an interdisciplinary pipeline where social scientists collaborate with NLP researchers across design, data, modeling, and evaluation. Social theory can guide scenario-based, context-rich prompts and treat annotator disagreement as a signal, while NLP methods scale these insights. Sustained collaboration replaces ad hoc work with iterative co-design and joint interpretation, yielding benchmarks that reflect fluid identities and models that engage culture as lived practice, improving both bias mitigation and interpretability.
By bridging critical social science with NLP practice, our framework not only critiques the status quo, but also offers a roadmap for more robust, inclusive, and dynamic cultural evaluation. It calls for models that do not simply conform to survey-driven metrics of alignment, but meaningfully engage with cultural difference as lived, negotiated, and plural, co-produced with the communities they are meant to serve.
@@
Our study focuses primarily on benchmarking methodologies rather than assessing the real-world impact of cultural benchmarks on LLM applications. Although improving benchmarks is a crucial step towards cultural sensitivity in AI, future research should investigate how these advancements translate to tangible improvements in AI-human interactions across various domains, including education, governance, healthcare and cross-cultural communication. Without real-world deployment analysis, the long-term effectiveness of improved cultural benchmarks remains an open question.
In addition, our position paper focuses on text-based evaluations, omitting vision-oriented culture benchmarks that interrogate the understanding of models of imagery, symbolism, and geospecific visual practices. Multimodal datasets such as CulturalVQA (visual question answering on food, clothing and rituals drawn from 11 countries) Nayak et al. (2024), CVQA Romero et al. (2024) (a culturally diverse and multilingual VQA set spanning 28 nations and 26 languages), and WorldCuisine Winata et al. (2025) (a dataset of world cuisines) demand analytical framework (e.g. visual-semantics grounding, region-specific iconography, and image–text alignment biases) that differ fundamentally from the linguistic constructs underlying our framework. Incorporating them would therefore require distinct taxonomies and fairness diagnostics, which we flag as an important avenue for future multimodal culture work.
We also recognize that we are not the first to introduce a taxonomy of cultural understanding in NLP Liu et al. (2025c). However, in contrast, our four-part taxonomy distinguishes between culture-as-knowledge (factual/symbolic information), culture-as-preference (shared values and moral attitudes), culture-as-dynamics (performative, enacted, context-dependent practices), and culture-as-bias (patterns of stereotyping and harm reduction). Our framework is not only descriptive, mapping how culture is conceptualized, but also prescriptive, challenging single-axis approaches, and recommending that future benchmarks intentionally combine these dimensions to better capture the complexity and contested nature of cultural life.
@@
Appendix A Case Study: CASA vs BLEND
This section contrasts the BLEND and CASA benchmarks, highlighting their distinct strengths in evaluating cultural understanding in large language models (LLMs).
BLEND excels in assessing culture-as-knowledge. It immerses models in realistic conversational scenarios that demand the contextual recall and appropriate application of culturally specific facts. For example, prompts might involve characters discussing preparations for Lunar New Year or Diwali, testing the model’s ability to refer to customary foods, greetings, or rituals in natural dialogue. This tests the integration of factual knowledge, such as understanding that joss paper is burned in Chinese ancestral rituals or that modaks are made during Ganesh Chaturthi.
In contrast, CASA is particularly strong in capturing culture-as-preference. It presents models with survey-style statements reflecting social or moral attitudes and assesses whether the models align their responses with the presumed dominant values of particular cultural groups. For example, CASA tasks might ask, "In your culture, is it acceptable to speak loudly in public spaces?" or assess whether a model expresses greater agreement with statements endorsing filial piety in East Asian contexts versus valuing individual autonomy in Western contexts.
However, both benchmarks operationalize culture in a certain way. BLEND’s interactive prompts assess cultural expertise by integrating object-level details that a model must integrate into natural discourse, reflecting authentic practice and symbolic reference. For example, BLEND might present a scenario in which characters plan a traditional celebration, such as discussing foods for the feast of Chuseok or customs for Ramadan, challenging the model to mention appropriate dishes (like exchanging songpyeon in Korea or breaking fast with dates in Muslim contexts). Meanwhile, CASA uses responses to attitude questions (such as the acceptability of public displays of emotion) to compare value patterns across cultural domains by mapping responses to expected normative judgments (like stoicism in Japanese settings versus expressiveness in Mediterranean cultures). Although BLEND challenges the breadth of LLM by introducing multiple details to recall for a culture, CASA may oversimplify by overlooking contestations or differences between cultures.
Appendix B Description of Reviewed Benchmarks